Consider a light ray incident upon a thin Plano-co
Answer #1

use the lens maker equation through the lens as

1/f = (n-1)* (1/R1 - 1/R2)

here R1 = 0.103 m

R2 for plane lens = infinity


so

1/f = (1.3 -1)*(1/0.103)

f = 0.343 m or 34.3 cm


so for this the answer is option 3, the ray will cross the axis 34.3 cm beynd the lens